OpenStudy (anonymous):

Find the product of the following by using methods of special products. 1. (2a^2+7)(2a^2-7) 2. (3x^2-yz^3)^2 3. (3x-5y)(5x+3y) 4. (3a^2b-7)^3 5. (2x-5)(4x^2-10x+25) Help again? Lol I'm so sorry i'm just not good at maaaaaath

OpenStudy (anonymous):

(2a^2+7)(2a^2-7) 2a^2 * 2a^2 = ___ 2a^2 * -7 = ____ 7 * 2a^2 = ___ 7 * -7 = ____ 

OpenStudy (anonymous):

just multiply

OpenStudy (anonymous):

why not give it a try @bottledthoughts

OpenStudy (jhannybean):

Or you could use the method of grouping. \[\large (2a^2+7)(2a^2-7)\]\[\large \color{green}{2a^2}(2a^2-7) +\color{green}7(2a^2-7)\]

OpenStudy (anonymous):

Or you could see that there's a special product called square difference that is (x+y)(x-y)=x^2-y^2 so, you just replace 2a^2=x and 7=y and voila...

OpenStudy (whpalmer4):

Yeah, difference of squares is the concept being requested here.

OpenStudy (whpalmer4):

(for the first and third ones)

OpenStudy (anonymous):

For number two you have a binomial square (x-y)^2=x^2-2xy+y^2 4 is a binomial cube (x-y)^3=x^3-3x^2y+3xy^2-y^3 And number five, factorize the second term and you will see magic happening (Remember about the perfect binomial square)